Divide 20 into two parts such that three times the square of one part exceeds the other part by 10
Answer by amarjeeth123(569)

Let the two parts be x and (20-x).
Then we have 3x^2-(20-x)=10
3x^2+x-30=0
3x^2+10x-9x-30=0
x(3x+10)-3(3x+10)=0
(x-3)(3x+10)=0
x=3 is the solution.
The two parts are 3 and 17 respectively.